OpenStudy (samigupta8):

cosπ/11+cos3π/11+cos5π/11+cos7π/11+cos9π/11

OpenStudy (anonymous):

your math teacher must hate youi

OpenStudy (samigupta8):

whu shud he i fact i hate him

OpenStudy (samigupta8):

plsssssssssss.... help me out

OpenStudy (anonymous):

i think you can set this equal to say \(x\) and then consider \(2x\sin(\frac{\pi}{11})\)

OpenStudy (samigupta8):

wht use will it make of...

Parth (parthkohli):

Hmm, my method does not work, bleh.

Parth (parthkohli):

What I posted wasn't even right.

OpenStudy (anonymous):

bunch of algebra to get \[2x\sin(\frac{\pi}{11})=\sin(\frac{10\pi}{11})\] and then \[2x\sin(\frac{\pi}{11})=\sin(\pi-\frac{\pi}{11})\[2x=1\] \[x=\frac{1}{2}\]

OpenStudy (anonymous):

let me see if i can find it for you written out, i can't write out all the algebra, hold on

OpenStudy (anonymous):

try this http://hk.knowledge.yahoo.com/question/question?qid=7008021104348